OpenStudy (anonymous):

t^3(4+i)(4-i)+sqrt-25

myininaya (myininaya):

Not sure what you want to do with this

myininaya (myininaya):

Let me know when you know. Thanks.

OpenStudy (anonymous):

simplify?

OpenStudy (anonymous):

yes

myininaya (myininaya):

\[t^4(4+i)(4-i)+\sqrt{-25}\] First of all I'm not really sure what simplify means here... \[\text{Secondly,} \sqrt{-1}=i\] \[\text{Thridly, } (a+i)(a-i)=a^2-i^2=a^2-(-1)=a^2+1\]

myininaya (myininaya):

And that is suppose to be t^3 not t^4

myininaya (myininaya):

So applying what I just said what does that give us so far?

myininaya (myininaya):

 Fill these blanks